OpenStudy (anonymous):

how do you solve x over 2 = 3x over 4 + 5

OpenStudy (anonymous):

\[\frac{x}{2}=\frac{3x}{4}+5\] \[4\times \frac{x}{2}=(\frac{3x}{4}+5)\times 4\] \[2x=3x+20\] \[-x=20\] \[x=-20\]